OpenStudy (anonymous):

The slope of ab is 5 over 3. Find the aquation of ab.

OpenStudy (anonymous):

someone please helpppp :)

OpenStudy (anonymous):

Your problem can not be solved as stated. You need the slope and a point on the line or two points on the line to get the equation of the line.

OpenStudy (anonymous):

okay maybe I got the slope wrong I will rewrite what it says in the book in full

OpenStudy (anonymous):

a (1, 4) and b( -2, -1) are two points. (i) Find the slope of ab (ii) Find the equation of ab The answer at the back of the book to (ii) 5x - 3y + 7= O but I want to know how you get this. The slope is definetely 5 over 3.

OpenStudy (anonymous):

Do you know the forumla for slope when given 2 points?

OpenStudy (anonymous):

no ?

OpenStudy (anonymous):

\[(y1-y2)/(x1-x2)\]

OpenStudy (anonymous):

That will give you your slope. Or "M".

OpenStudy (anonymous):

Now that you have your slope pick one of the points and apply it to the point slope forumla. y-y1=m(x-x1)

OpenStudy (anonymous):

y1 and x1 are points on your line..... (1,4) or (-2,-1)

OpenStudy (anonymous):

okay thank you so much you totally helped me out :)